# DeepTx: Real-Time Transaction Risk Analysis via Multi-Modal Features and LLM Reasoning

### v1

@@ I Introduction
Several tools have been developed to mitigate phishing transactions, each adopting different detection strategies with varying levels of openness and effectiveness. PTXPhish [1] is an open-source tool that analyzes behavioral and contextual features within transactions. However, its rule-based design lacks adaptability and often fails to detect previously unseen attack vectors. In contrast, ScamSniffer [3] and Pocket Universe [4] focus on call data simulation, matching results against hard-coded patterns. While ScamSniffer partially discloses its block lists and browser extension code, both tools retain proprietary detection logic, limiting their transparency and extensibility. Forta [5], a decentralized alerting system, employs community-operated bots to monitor live on-chain activity. Although Forta supports real-time data streams, its alerts are typically issued only after transactions have been mined, and together with other tools that rely on static rules or post-hoc pattern recognition, it offers limited protection against phishing tactics involving UI forgery, calldata obfuscation, or execution flows at the time of signing.
In this paper, we propose DeepTx, a tool that simulates user-intended transactions as if they were already signed, and extracts multi-modal features including behavioral traces, contextual behaviors, and user interface indicators. It then uses LLMs to analyze the transaction’s intent and potential security risks. The system generates a natural language risk report that includes a severity level, explanatory description, and suggested user action. To enhance reliability, DeepTx performs consensus checking across multiple LLMs and applies self-reflection techniques to resolve inconsistencies or ambiguous results.
@@ II-E3 Fallback Voting Mechanism
If no consensus is reached after $M$ rounds of self-reflection, DeepTx falls back to a weighted voting strategy, as shown in the second part of Algorithm 1. The final risk label is selected based on the cumulative confidence-weighted votes for each class. Specifically, the total score for each candidate label $r$ is computed as:
$\text{Score}(r)=\sum_{i=1}^{n}\mathbb{I}[r_{i}=r]\cdot c_{i}$
where $r_{i}$ is the label assigned by model $i$ with associated confidence $c_{i}$, and $\mathbb{I}[\cdot]$ is the indicator function. The final predicted label $r^{*}$ is selected as:
$r^{*}=\arg\max_{r}\text{Score}(r).$
The output from any model satisfying $r_{i}=r^{*}$ is used as the final decision. The system merges all available outputs from the final round to produce a user-facing summary that includes supporting explanations, risk rationale, and suggested actions.
@@ II-F Output Module
III Tool Usage
DeepTx can be used after installing Python dependencies and setting environment variables such as RPC_URL, ETHERSCAN_API_KEY, and OPENAI_KEY. Given a transaction prepared on the RPC node, run python3 main.py <transaction_hash> to retrieve and analyze it. The tool returns a structured report describing its semantic properties.
IV Preliminary Evaluation
@@
IV-A Dataset Collection
We manually labeled 12 phishing cases from two sources, including 5 simulation-based phishing challenges from UnPhishable [9] and 7 real-world cases collected from ScamSniffer database and archived phishing websites [12]. These cases cover a range of semantic and deception strategies, including malicious token approvals, proxy spoofing, and impersonation via fake interfaces. For each transaction, we collected (1) the associated frontend webpage and JavaScript if available, (2) verified or decompiled smart contract code using Heimdall, and (3) the original victim transaction from the chain.
@@ IV-A Dataset Collection
IV-B Experiment Setup and Metrics

### v2

@@ I Introduction
Several tools have been developed to mitigate phishing transactions, each adopting different detection strategies with varying levels of openness and effectiveness. PTXPhish [2] is an open-source tool that analyzes behavioral and contextual features within transactions. However, its rule-based design lacks adaptability and often fails to detect previously unseen attack vectors. In contrast, ScamSniffer [3] and Pocket Universe [4] focus on call data simulation, matching results against hard-coded patterns. While ScamSniffer partially discloses its block lists and browser extension code, both tools retain proprietary detection logic, limiting their transparency and extensibility. Forta [5], a decentralized alerting system, employs community-operated bots to monitor live on-chain activity. Although Forta supports real-time data streams, its alerts are typically issued only after transactions have been mined, and together with other tools that rely on static rules or post-hoc pattern recognition, it offers limited protection against phishing tactics involving UI forgery, calldata obfuscation, or execution flows at the time of signing. An early LLM-based attempt at transaction analysis exists, but it lacks a clear architecture and an end-to-end implementation [6].
In this paper, we propose DeepTx, a tool that simulates user-intended transactions as if they were already signed, and extracts multi-modal features including behavioral traces, contextual behaviors, and user interface indicators. It then uses LLMs to analyze the transaction’s intent and potential security risks. The system generates a natural language risk report that includes a severity level, explanatory description, and suggested user action. To enhance reliability, DeepTx performs consensus checking across multiple LLMs and applies self-reflection techniques to resolve ambiguous results.
@@ II-E3 Fallback Voting Mechanism
If no consensus is reached after $M$ rounds of self-reflection, DeepTx falls back to a weighted voting strategy, as shown in the second part of Algorithm 1. The final risk label is selected based on the cumulative confidence-weighted votes for each class. Specifically, the total score for each candidate label $r$ is computed as: $\text{Score}(r)=\sum_{i=1}^{n}\mathbb{I}[r_{i}=r]\cdot c_{i}$ where $r_{i}$ is the label assigned by model $i$ with associated confidence $c_{i}$, and $\mathbb{I}[\cdot]$ is the indicator function. The final predicted label $r^{*}$ is selected as: $r^{*}=\arg\max_{r}\text{Score}(r).$ The output from any model satisfying $r_{i}=r^{*}$ is used as the final decision. The system merges all available outputs from the final round to produce a user-facing summary that includes supporting explanations, risk rationale, and suggested actions.
@@ II-F Output Module
II-G Tool Usage
After configuring the .env file, DeepTx provides two modes—Historical Transaction and Simulation—to analyze existing on-chain transactions or execute a specified contract function for simulated execution, returning a structured semantic report in both cases.
III Preliminary Evaluation
@@
III-A Dataset Collection
We manually labeled 12 phishing cases from two sources, including 5 simulation-based phishing challenges from UnPhishable [UnPhishable] and 7 real-world cases collected from ScamSniffer database and archived phishing websites [8]. These cases cover a range of semantic and deception strategies, including malicious token approvals, proxy spoofing, and impersonation via fake interfaces. For each transaction, we collected (1) the associated frontend webpage and JavaScript if available, (2) verified or decompiled smart contract code, and (3) the original victim transaction from the chain.
@@ III-A Dataset Collection
III-B Experiment Setup and Metrics